# Prompting LLMs for Code Editing: Struggles and Remedies

### v1

@@ 1. Introduction
First, in Section 3, we carry out, to our knowledge, the first log-based study of developers’ behavior when it comes to prompting for code transformations. Previous studies in the area of user experience with AI-based tools have focused on asking developers how they feel through surveys or interview studies (D’Angelo et al., 2024), or through exploratory user studies (Vaithilingam et al., 2022; Barke et al., 2023) that try to elicit user opinions on the usefulness of AI-based coding tools and on the trust developers place in them. Since we have extensive fine-grained logs of developer interactions with this AI-based feature in a widely used IDE at a large company, we have the opportunity to make behavioral observations. For instance, we can tell how often, and to what extent, developers tend to undo and repeat the exact same prompt, and to what extent they accept an AI-based suggested code edit only to heavily change it afterwards.
Second, we approximate the “ground truth” answer expected by the developer, by looking at a future code snapshot or developer-submitted feedback for the code suggestion. This allows us to as ask, counterfactually, what would have been a prompt that produced that ground truth answer, or something very close to it. From a semi-automated study that enhances user prompts by given the ground truth edits, we find that 59.72% of the unsatisfactory code edits could be improved to be semantically equivalent to the ground truth. We also conduct a qualitative study to understand the commonly missing or under-specified details in the developer prompts by analyzing the differences between the original and the improved prompts. We find five common types of information missing from prompts: specifics, operationalization plan, localization/scope, codebase context, and developer intent.
Finally, rather than ask developers to work harder, we propose tool support for automatic prompt enhancement in Section 6. Here, the desired code edit is not known, but based on the categories of commonly missing or under-specified elements in developer prompts from the above study, we automatically augment—using an LLM—a prompt to append it with (likely) relevant contextual information to it. We find that this tool—called AutoPrompter—can improve the effectiveness of developer prompts – which are generally lacking in specificity and context – by around 27%, automatically and without further developer input.
Motivating example. Table 1 shows—for the same example as used in Figure 1—the original developer prompt, a prompt improved by AutoPrompter, the code edits generated by Transform Code using the two prompts, as well as the desired code edit. The original developer prompt, “simplify this” neither clearly indicates the developer’s goal, which is to simplify the code to be more concise (i.e., user intent), nor specifies how to simplify the code using Python’s built-in list comprehension (i.e., operationalization plan), based on the desired code edit. This led Transform Code to originally produce simplified code, but by changing the behavior of the code. By identifying these missing specifications and reasoning based on the code context provided, AutoPrompter produced a prompt that makes the code more concise using list comprehension, while keeping the functionality intact.
Summary of results.
  • Around 7% of the Transform Code requests are identical re-prompting requests; moreover, even when developers edit their prompts at the same location, many do not make significant changes. This indicates that developers are reluctant to invest too much work in revising their prompts in the context of Transform Code. [See Section 3]
  • Our qualitative analysis constructs a codebook of commonly missing information from developers’ prompts: not being specific, missing an operationalization plan, faulty localization/scope, missing code base context or intent. This is done by automatically extracting ground-truth edit information from historical logs and code repository state. [See Section 4 and Section 5]
  • It is possible to automatically enhance the developer’s prompt for Transform Code without further manual input: about 27% of the unsatisfactory code edits in the test set were improved by automatically updating the prompt with AutoPrompter. [See Section 6]
2. Background & Related work
2.1. AI Developer Tools for Professional Developers
In recent years, driven by significant improvements in large language models (LLMs), many developer tools have been built upon or have incorporated LLMs. GitHub Copilot (cop, 2025) is one of the earliest LLM-powered developer tools, suggesting code in real time, based on context. More similar tools have then built, such as Cursor (cur, 2025), Windsurf (win, 2025), and Bolt (bol, 2025), supporting code completion, editing, and generation, or even agent-based support. Significant effort has also been invested in AI-enhanced software development tools for internal use, such as at Meta (Bader et al., 2021) and Google (Chandra and Tabachnyk, 2024; Tabachnyk, 2022; Froemmgen et al., 2024), to support their developers’ workflows and proprietary codebases. Many have conducted studies to understand their potential to enhance developer productivity, investigating both perceived (Bird et al., 2023; Murillo et al., 2023) and actual (Peng et al., 2023; Paradis et al., 2024) productivity gains, through various methodologies. Research has also studied challenges developers face when using AI tools. A notable struggle reported multiple times is in effectively communicating requirements to AI tools, which often requires developer cognitive effort (Parameswaran et al., 2024; Nam et al., 2024; Vaithilingam et al., 2022; Liu et al., 2023; Tafreshipour et al., 2024).
Analyzing telemetry logs to understand usage patterns (Murphy-Hill et al., 2018; Nam et al., 2024) and improve features (Bibaev et al., 2022; Froemmgen et al., 2024) is a common practice in software engineering. However, AI tools have only recently been broadly adopted by developers for real-world usage and are rapidly evolving. Consequently, research on how developers incorporate these tools into their actual workflows has been limited. More recently, some efforts have been made to measure the overall impact of AI tools at scale, such as by measuring acceptance rates (Ziegler et al., 2022) or the number of pull requests (Peng et al., 2023). However, these studies primarily focused on quantifying productivity gains rather than understanding how developers use the tools. One notable exception is Mozannar et al. (2024), which collected and analyzed telemetry data from programmers using Copilot. While that paper provides valuable insights into programmer activities using telemetry segments, the data was collected from only 21 participants and the analysis was done after manual labeling. In contrast, our analysis examines logs from developers at Google, reporting insights at a large scale.
2.2. Existing Approaches for Prompt Enhancement
As early research has identified challenges in writing effective prompts (Parameswaran et al., 2024; Nam et al., 2024; Vaithilingam et al., 2022; Liu et al., 2023), researchers and practitioners have explored ways to assist developers in prompt writing. Several (semi-) automatic approaches have been proposed, e.g., meta-prompting (Yang et al., 2024; Fernando et al., 2024; Zhou et al., 2023) for automatically optimizing prompts in a systematic fashion. As another example, it was demonstrated in Automatic Prompt Engineer (APE) (Zhou et al., 2023) that LLMs can discover human-level prompts through a search process, outperforming manually written prompts on multiple benchmarks. However, the main research efforts have focused on optimizing the system instruction prompt that can be used for several distinct tasks within the same category, often requiring many examples to automatically quantify the quality gains from prompt improvement. Another notable example are prompt augmentation tools for image generation, like LMOps (Hao et al., 2023), which aim at generating diverse images through augmented developer prompts. However, their main goal is to generate various outputs, rather than satisfying more constraints by inferring missing specifications from the surrounding context.
3. Transform Code Log Analysis
Transform Code has been released to Google’s developers for about a year, and has since been adopted by a large fraction of developers, who have varying degrees of expertise with ML. Through iterative improvements to both the underlying LLM and the user interface, Transform Code has become a core part of developers’ workflows. Despite its positive reception and appreciation, generating code edits from natural language prompts remains a relatively new concept for developers. Thus, some developers have expressed challenges, mirroring the experiences of users of similar tools (Kruse et al., 2024; Nam et al., 2024).
In the ideal case, a developer using Transform Code would send a request with a prompt for a given task, Transform Code would provide an appropriate code edit, the developer would review the edit and proceed to the next task. However, in practice, it is unrealistic to expect an LLM-powered tool to provide the perfect solution on the first attempt. In this section, we analyze telemetry logs from Transform Code and the IDE, focusing on how developers use Transform Code, and how they react when Transform Code does not provide the ideal code edit they anticipated.
3.1. Data
3.1.1. In-IDE Event Logs
3.1.2. Transform Code Logs
3.1.3. Dataset used for analysis
From this combined dataset, we sampled data of individual developers who used Transform Code at least 24 times during the month, which is the median usage of the entire population. This threshold was set to minimize noise arising from non-regular usage of Transform Code, e.g., where new users might test Transform Code with simple prompts like ”hi.”
3.1.4. Limitations
(a) After rejection.
(b) After acceptance.
3.2. Post-Transform Code interactions
Figure 2 visualizes the two consecutive in-IDE developer actions after requesting Transform Code. We first analyzed the actions made after rejection, to understand how developers cope with unsatisfactory Transform Code edits. Then, we also analyzed the actions made after acceptance, Figure 2-(a), to check for any differences in developers’ reactions. This distinction is also important because the Transform Code edit is applied to the user’s code upon acceptance, but discarded upon rejection. Thus, subsequent edits made by the user in the same parts of the code after acceptance likely indicate refinement or adjustment of the applied Transform Code edit, whereas actions taken after rejection indicate the user is continuing to work on the original code.
3.2.1. Post-rejection
The most common subsequent action after rejecting a Transform Code edit was Type (33.5%), including typing a character or inserting a line. This indicates that developers often start working on their tasks by themselves after checking an unsatisfactory Transform Code edit. A particularly interesting next action was a repeated request to Transform Code (24.3%), which is the second most common action after rejecting a Transform Code edit. This likely indicates that developers re-prompted Transform Code for the same task. We found that not an insignificant proportion of developers (22.2%) made even another Transform Code request, possibly indicating developer struggle in using Transform Code, and/or their desire to delegate the task to Transform Code instead of picking up on the task themselves, like those who started typing. The following frequent subsequent actions were Delete (20.1%), Paste (9.2%) and Cut (3.1%), showing other mechanisms of editing the original code, possibly adopting code from code or documentation search.
3.2.2. Post-acceptance
The most common subsequent action was Delete (32.9%) and the second most frequent action was Type (28.7%). This indicates that the developers frequently make edits to the Transform Code-edited code, after accepting a response, to further refine it. The next most frequent action was a repeated request to Transform Code (11.9%). Even after accepting the Transform Code edit, users requested Transform Code from the same code region more than 10% of the time. This could indicate that developers found unsatisfactory edits after accepting the edits, and re-prompted for the same task, or it is still possible that developers moved on to the next task and requested another code edit within the same part (e.g., adding documentation after implementing a function). Developers also made edits via other mechanisms, possibly to the Transform Code-edited code, by pasting code (5.8%), or cutting the code (5.1%).
3.3. Characteristics of Re-prompting
3.3.1. Frequency of Identical Re-prompting
We first investigated how often developers re-prompt with an identical prompt. We started with identical re-prompting, as it is very likely that the user is making a follow-up request, instead of moving on the next task after getting a satisfactory response.
3.3.2. Re-prompting with tweaks
3.4. Discussion
Overall, we observed that developers often do not invest sufficient time in crafting effective prompts when using Transform Code. Although documentation and the Transform Code user interface instructed users to provide specific details and necessary context, we found that few users did so when tweaking the prompt. Developers typically resorted to manual editing or invested minimal effort in refining the prompt for a subsequent attempt when an initial, low-effort attempt with Transform Code yielded unsatisfactory code edits. This behavior suggests that there may be an ‘economics of prompting,’ where developers weigh the costs of prompting against the expected benefits. Thus, to improve the user experience and maximize the benefits of Transform Code, enhancing the underlying model performance alone is not sufficient; efforts must also focus on reducing users’ cognitive load during prompting and on helping them write prompts more effectively.
4. Real-world Code Editing Data
To inform the development of solutions that help developers write more effective prompts for Transform Code, our initial objective was to understand current prompting practices and identify areas for improvement. Towards this goal, we collected a dataset of real-world Transform Code usage, specifically focusing on developer requests that yielded not satisfactory outcomes.
4.1. User Submitted Desired Code Edits
Transform Code has a feedback mechanism, where developers unhappy with the code edit suggestion from Transform Code can submit their desired code edit instead. This improved suggestion is bundled with the original prompt, code context, and original Transform Code suggested code edit upon submission. For these examples, the desired code edits most probably satisfy the user intent and expectations, as developers themselves need to deliberately create and submit the improved rewrite. However, not all developers are motivated to invest time in submitting such rewrites; therefore, the number of these examples are quite limited. After excluding the examples used for the fine-tuning of the Transform Code model, we had 52 examples left for this analysis.
4.2. Automatic Extraction of Desired Code Edits
Figure 4. Developers make changes to the code (see Section 3.2) to fully resolve the task described in the prompt for the request, regardless of how (un)satisfied they are with the Transform Code edits are. To obtain the ground-truth code edit given the user request to Transform Code, we extract a future snapshot of the code after the makes these edits.
To obtain the desired code edit, we extracted a future snapshot of the code, assuming that regardless of how (un)satisfied the developers were, the task would have been resolved at some later point in time (see Figure 4). We consider the code state after 30 minutes, to leave enough time for developers to make their desired edits, especially for cases where the task requires finding additional information. However, given that the edit time can vary significantly, depending on the task complexity and the user expertise, it was possible that code edits not relevant to the original Transform Code request could have been included. Developers could also potentially give up on finishing the task, or work on different tasks in the meantime. This made manual verification necessary.
  • the desired code edit only removes code from the original code.
4.3. Full Dataset for Analysis and Solution Building
Combining the data collected from the two sources resulted in 159 examples. Among these, we used 111 examples (70%) of the dataset for understanding developer prompting practices and issues, and we set aside 48 (30%) to be used for testing the solution we would build based on our gained understanding. Among these, we excluded examples that request code documentation only or simple renaming, where it is challenging to understand why the original Transform Code code edits were not satisfactory, and there is a subjective component involved (see Table E-Example 3 in Appendix). This resulted in a training dataset of 72 examples, and a test dataset of 33 examples. As the data were collected from real-world logs, we cannot share the entire dataset as it contains proprietary code, but we can only report the results and share examples.
5. Understanding Common Issues of Developer Prompts
With the collected real-world dataset, we conducted an error analysis of the developer requests that resulted in unsatisfactory code edits from Transform Code, to understand what would be required in a “better prompt”. We leveraged an LLM111in our case, Gemini, but we could have used any of the popular LLMs. to help us identify–for each example–the missing ingredients in a prompt, given prior knowledge of the complete code context, the prompt and the desired edit. This is based on a premise that with such a “better prompt”, Transform Code would likely have generated the desired solution.
5.1. Error Analysis using an LLM
To collect the common issues with user prompts (i.e., what is missing from the prompt so that Transform Code gives the desired result), we presented the LLM with: (1) the original user prompt, (2) the code context, (3) the original Transform Code edit, and (4) the desired code edit. Then we asked the LLM to produce (a) an explanation of what was missing from the original user prompt and (b) an improved prompt. See Appendix A for the full system prompt used.
5.1.1. Feasibility Testing
As a sanity check to test whether an LLM can reason about the issues in the user prompts, we tested whether Transform Code can replicate the desired edits from the LLM-enhanced prompts, when provided with the desired edits to enhance the prompt. Out of 72 examples, we found that 43 enhanced prompts (59.72%) enabled Transform Code to replicate the desired edits. These numbers indicate that the LLM is able to pinpoint the issues with the user prompts well, and with all the provided information in the prompt, Transform Code is able to generate satisfactory outcomes for a fairly large number of cases. At the same time, we found that not every unsatisfactory Transform Code edit is due to an incomplete user prompt. For example, cases that cannot be easily abstracted into a concise natural language prompt (e.g., adding tens of import files without clear patterns across them) could not be improved by rewriting prompts even after seeing the desired edit. We also discovered that in some examples, Transform Code provided better code edits than the “desired code edits.” Here the developer may have accepted the Transform Code suggestion, had they seen it, but as we do not know the developer’s original intent, they were not considered semantically identical. Finally, the underlying model has limitations; it may not be able to generate the desired code edit even when presented with the perfect prompt.
5.2. Qualitative Analysis
We qualitatively analyzed the issues of user prompts for Transform Code extracted from the 72 examples in our dataset, using thematic analysis (Braun and Clarke, 2006). Our qualitative analysis consisted of several phases. Three authors analyzed the original user prompts and the feedback provided by the LLM, which indicated what information was missing from the original prompt using inductive coding (Fereday and Muir-Cochrane, 2006). As we reviewed the data, we generated an initial set of categories to describe the different types of missing context and specifications. If we could not clearly identify which information was missing from the original prompt, we referred to the code context, the original Transform Code edit, and the desired code edit, which also helped mitigate any errors coming from potential hallucinations. We also wrote analytic memos during the discussion, including observations of patterns and potential connections among the categories.
5.3. Results
Code # Description
Specifics 28 Missing or under-specified specifications, for example the prompt lacks value information
Operationalization plan 28 Missing or under-specified information on what operations need to be performed or what functions need to be used, or what edge case handling needed to get to the desired code edit
Codebase context 11 Prompt itself did not provide necessary background information for the model to be able to generate the desired code edits, including missing information about the broader system, existing codebase. Developers may have assumed that the model had all the necessary context.
User intent 12 Missing or under-specified user intent, in the prompt it is not clear what the developer’s goal is or what type of code edit they want to achieve
The analysis resulted in 5 codes of missing or under-specified information in user prompts for Transform Code. Table 2 shows the codebook (Ando et al., 2014), including the code, the frequency of the codes, and the description. Each example could be labeled with up to 2 codes.
Codebase Context. Codebase context refers to cases where the user assumes that Transform Code would know all the background knowledge and details about the codebase they are working on, in cases where Transform Code actually had no access to such information. This could happen because end users might not have a good understanding of how the underlying model is trained, which data are used, and what additional context us included as part of the request, which may have led the developers to assume a level of contextual awareness that the model does not possess. For example, the desired code edit that Transform Code was not able to produce given “add [library]” was successfully replicated with an improved prompt, “Add ”[path to the library]” as a dependency to the [build] rule.
User Intent. A frequent example of this category is a user asking Transform Code to perform a rather vague and open-ended task. For example, for a request with the original user prompt “make it simpler”, the LLM pointed out that “The prompt did not specify the intended behavior when [variable1] is present in [variable2]. The model incorrectly interpreted “make it simpler” as removing the redundant assignment instead of preserving the original conditional logic with the necessary ‘continue‘ statement.” The improved prompt generated by the LLM, “Refactor the conditional logic for assigning [variable3] to maintain the original behavior, adding a ’continue’ statement when [variable1] is in [variable2].”, successfully replicated the desired code edit. When prompts lack user intent, compared to other issues, it becomes extremely difficult for Transform Code to produce the desired code edit, as it almost requires reading the minds of developers. Thus, if the desired code edit matches the most likely code edit to be made in general, given the code context, Transform Code can produce a good code edit, but if not, it is very challenging to achieve the desired code edit.
Satisfactory Prompts. We identified 5 examples from the dataset where the original user prompts appeared satisfactory without requiring further improvement, at least from the researchers’ perspective. Because the dataset is a collection of user requests that resulted in unsatisfactory code edits, rather than a collection of inadequate prompts, it is reasonable that some of these unsatisfactory code edits are not attributable to the user prompts.
5.4. Feasibility of Automatic Analysis
In this study, we conducted a qualitative analysis following the traditional thematic analysis process, as this was the first attempt to understand user prompt issues for LLMs in the context of code editing. However, this process of extracting insights from usage data could potentially be automated with LLMs, similar to other attempts in conducting qualitative research with LLMs (Bano et al., 2023; Meng et al., 2024), particularly to reduce researcher effort and cost. To assess the feasibility of such automation, we instructed an LLM to carry out the open coding and codebook construction. We found the LLM-generated codebook to be largely similar to the one created by researchers (see Appendix C). This result suggests the potential for applying automated analysis with low cost to understand user challenges in other tool contexts (e.g., automatic program repair (Rondon et al., 2025)) or domains (e.g., code migration (Nikolov et al., 2025)).
6. Automatic User Prompt Enhancement
In the previous section, we identified categories of common issues with user prompts. In this section, we tested the feasibility of automatically improving user prompts by exploiting this finding, and tested if some of the missing or under-specified user requirements could be inferred from the code context.
@@ 6.1. Approach to Prompt Enhancement
To test the feasibility of automatically enhancing developer prompts by inferring the developer intent solely based on the developer prompt and to determine how much automatic prompt enhancement can help Transform Code generate the desired code edit, we built and evaluated AutoPrompter. Given (1) the original developer prompt, (2) Transform Code suggested edit generated with the original developer prompt, and (3) the code context provided by the developer, AutoPrompter was instructed to check whether the given developer prompt misses or under-specifies any of the key information, and to infer them from the given code context to enhance the prompt. We also instructed AutoPrompter to declare if it cannot confidently infer an improved prompt because of a lack of information, and ti explicitly request additional context instead of giving the improved prompt.
As the LLM is inferring the missing specifications from the code context, it is very likely that there can be multiple probable ways of further specifying the developer prompt. For example, when a developer asks to “document”, it can be “improve documentation to be more concise” (user intent) but also “write docstrings for every function” (localization/scope). To make sure that AutoPrompter generates diverse prompts that can cover multiple possible specification options, we instructed it to provide missing/unspecified instructions given the categories we found in the previous section as a reference, and to improve prompts considering each of those categories. See Appendix D for the full system prompt used for AutoPrompter.
@@ 6.2.2. Metrics
ChrF (Character-level F-score) (Popovic, 2015). Computes an $F_{1}$ score based on the proportion of character-level $n$-grams.
@@
Autorater (Zheng et al., 2023). A complement to text-based similarity scores, that can suffer from syntactic differences; measures the semantic similarity between code edits using an LLM-as-a-judge. Given the original developer prompt, the AutoPrompter enhanced prompt, and the desired code edits, the LLM evaluates the similarity; i.e., how similar the code edits are to the desired code edit.
Human evaluation. Two authors separately evaluated the improved code edits by comparing them to the ground-truths and the original user prompts. Each example was labeled as 1 when the improved code edit was semantically equivalent to the ground truth considering the original user prompt and the code context, and 0 if not. After separate evaluation, we found a substantial agreement by achieving a 0.86 Cohen’s Kappa score. For cases where the two authors did not agree, we averaged the two scores.
@@
The other popular text based metrics, such as BLEU (Papineni et al., 2002), are not suitable for assessing the quality of code edits. We also considered measuring the CodeBLEU (Ren et al., 2020) scores for the evaluation. However, as CodeBLEU is language specific in implementation and given the variety of languages our benchmark contains, we decided that extending the CodeBLEU implementation was not practical.
@@
6.2.3. Limit Study
6.3. Results
Table 3 summarizes the results. In interpreting the data, we focused on human evaluation results, as we found that text-based similarity scores often do not correlate with human evaluation, especially when the differences are small. Moreover, text-based metrics often give unsatisfactory results in the case of syntactic differences that do not matter semantically.
The semantic human evaluation count is 0 by design, as the dataset was built to only include unsatisfactory Transform Code code edits, i.e., semantically different from the desired code edit.
Recall from Section 6.1 that we ask the LLM to produce multiple prompt variants. Given multiple prompts, each one improved with a different strategy, we used different ways of selecting the best improved prompt to use. In Table 3, self-selection, farthest, and llm show the performance of different selection approaches. First, for (self-selection), we instructed the AutoPrompter directly to select a prompt that is most likely to elicit the desired code edit. We also tried choosing a prompt for which Transform Code produces the most distinctive code edit from the original code edit (farthest). This is based on the idea that if the developer was not satisfied with the original code edit, the desired code edit will more likely be found within the code edit candidates that are different from the unsatisfactory code edit, than within those similar to the unsatisfactory one. Finally, we used an LLM to select the best improved Transform Code code edit, given the original developer prompt (llm).
Overall, the evaluation demonstrates the feasibility of automatically improving the developer prompt: 9 out of 33 (27%) examples were improved by AutoPrompter, which indicates that it is often feasible to infer missing specifications from the given code context. This seems like a good result especially when considering 37% as the limit (as computed by the process in  Section 6.2.3) for this test set.
Recall also from Section 6.2.3 that some of the examples may not be amenable to this process of prompt enhancement; and, finally, Transform Code also has limits in terms of the power of the model itself in interpreting the prompt and carrying out a desired edit.

### v2

@@ 6.1. Approach to Prompt Enhancement
To test the feasibility of, and the extent to which, automatically enhancing developer prompts, we built and evaluated AutoPrompter using an LLM (see Figure 4). Given (1) the original developer prompt, (2) Transform Code suggested edit generated with the original developer prompt, and (3) the code context provided by the developer, AutoPrompter was instructed to check whether the given developer prompt misses or under-specifies any of the key information, and to infer them from the given code context to enhance the prompt. We also instructed AutoPrompter to declare if it cannot confidently infer an improved prompt because of a lack of information, and to explicitly request additional context instead of giving the improved prompt.
As the LLM is inferring the missing specifications from the code context, it is very likely that there can be multiple probable ways of further specifying the developer prompt. For example, when a developer asks to “document”, it can be “improve documentation to be more concise” (user intent) but also “write docstrings for every function” (localization/scope). To make sure that AutoPrompter generates diverse prompts that can cover multiple possible specification options, we instructed it to provide missing/unspecified instructions given the categories we found in the previous section as a reference, and to improve prompts considering each of those categories. The prompt used can be found in Appendix D (artifact).
@@ 6.2.2. Metrics
ChrF (Character-level F-score) (popovic2015chrf). Computes an $F_{1}$ score based on the proportion of character-level $n$-grams.
@@
Autorater (zheng2023judging). A complement to text-based similarity scores, which can suffer from syntactic differences, measures the semantic similarity between code edits using an LLM-as-a-judge. Given the original developer prompt, the AutoPrompter enhanced prompt, and the desired code edits, the LLM evaluates the similarity; i.e., how similar the code edits are to the desired code edit.
Human evaluation. Two authors separately evaluated the improved code edits by comparing them to the ground-truths and the original user prompts. Each example was labeled as 1 when the improved code edit was semantically equivalent to the ground truth, considering the original user prompt and the code context, and 0 if not. After a separate evaluation, we found a substantial agreement by achieving a 0.68 Cohen’s Kappa score. For cases where the two authors did not agree, we averaged the two scores.
@@
The other popular text-based metrics, such as BLEU (papineni2002bleu), are not suitable for assessing the quality of code edits. We also considered measuring the CodeBLEU (ren2020codebleu) scores for the evaluation. However, as CodeBLEU is language-specific in implementation and given the variety of languages our benchmark contains, we decided that extending the CodeBLEU implementation was not practical.
6.2.3. Limit Study
6.3. Results
@@
Table 3 summarizes the results. In interpreting the data, we focused on human evaluation results, as we found that text-based similarity scores often do not correlate with human evaluation, especially when the differences are small. Moreover, text-based metrics often give unsatisfactory results in the case of syntactic differences that do not matter semantically. The semantic human evaluation count is 0 by design, as the dataset was built to only include unsatisfactory Transform Code code edits, i.e., semantically different from the desired code edit.
As two baselines, we used original and none. original displays the scores of original Transform Code edits, based solely on the original user prompts. none is based on the scores of Transform Code response generated only with the LLM prompt enhancement, but without including Table 1 as part of the instruction. The difference between this and other scores thus indicates the usefulness of instructing AutoPrompter to consider the common issues of user prompting. We have multiple versions of AutoPrompter, where the simplest is single, which asks the LLM to produce one enhanced prompt given the information. Also, recall from Section 6.1 that we ask the LLM to produce multiple prompt variants. Given multiple prompts, each one improved with a different strategy, we employed various methods to select the best improved prompt to use. In Table 3, self-selection, farthest, and llm show the performance of different selection approaches. First, for (self-selection), we instructed the AutoPrompter directly to select a prompt that is most likely to elicit the desired code edit. We also tried choosing a prompt for which Transform Code produces the most distinctive code edit from the original code edit (farthest). This is based on the idea that if the developer is not satisfied with the original code edit, the desired code edit is more likely to be found within code edit candidates that differ from the unsatisfactory code edit than within those similar to it. Finally, we used an LLM to select the best improved Transform Code code edit, given the original developer prompt (llm).
Overall, the evaluation shows the feasibility of automatically improving the developer prompt: 9 out of 33 (27%) examples were improved by AutoPrompter, which indicates that it is often feasible to infer missing specifications from the given code context. This seems like a good result, especially when considering 37% as the limit (as computed by the process in  Section 6.2.3) for this test set. Among the AutoPrompter variations, self-selection appears to perform the worst; this could indicate that it might not be easy for an LLM to tell whether the given prompt meets the user intent or not, without reviewing the code edits generated by it. Recall also from Section 6.2.3 that some of the examples may not be amenable to this process of prompt enhancement; and, finally, Transform Code also has limits in terms of the power of the model itself in interpreting the prompt and carrying out a desired edit.